# Lower semicontinuity, Stoilow factorization and principal maps

### v2

@@ 1.1. Principal maps
Notice that if $f\in\mathrm{C}^{1}(\mathbb{D})\cap\mathrm{W}^{1,1}(\mathbb{D})$ with $\det\textup{D}f>0$, and for some linear $A\in\mathbb{R}^{2\times 2}_{+}$ we have $f-A\in\mathrm{W}^{1,1}_{0}(\mathbb{D})$, so $f$ is a classical test function for the quasiconvexity inequality, then $f$ can be extended as principal map off the unit disc. For example if $A(z)=b_{0}z+b_{1}\bar{z}$, then the map $f^{\rm prin}(z)=f(z)\mathbf{1}_{\mathbb{D}}+\mathbf{1}_{\mathbb{C}\setminus\mathbb{D}}\bigl{(}b_{0}z+\frac{b_{1}}{z}\bigr{)}$ is principal in the sense of Definition 1.1. Therefore principal quasiconvexity implies a standard quasiconvexity inequality for sufficiently smooth maps. However, from the classical area formula we that $-\det(A)$ is principal quasiconvex, while the determinant itself is not. That is, quasiconvexity does not imply principal quasiconvexity in general.
@@
On the other hand, the class of boundary values of all principal maps is much larger than just the linear ones. In fact, up to a Möbius transform, the class consists of the traces on $\partial\mathbb{D}$ of conformal maps of $\mathbb{C}\setminus\overline{\mathbb{D}}$ onto Jordan domains $\Omega$, with $\infty\in\Omega$ and boundary $\partial\Omega$ sufficiently regular.
@@
Let us then briefly explain why principal quasiconvexity is relevant in dealing with variational integrals. The proof of Morrey [38] for the lower semicontinuity of functionals with standard quasiconvex integrands goes basically as follows. Firstly, given a weakly converging sequence $\{\textup{D}\psi_{j}\}$ of Sobolev functions, one localises the sequence to achieve convergence to a linear map. Second, one replaces this localized sequence with one of affine boundary values, $u_{j}-Az\in\mathrm{W}_{0}^{1,2}(\Omega)$, via a carefully chosen cutoff. And third, one applies directly the definition of quasiconvexity to each such $u_{j}$.
However, such a strategy breaks down for functionals defined on +$\mathbb{R}^{2\times 2}_{+}$ because the cut-off modification might give negative determinants on sets of positive measure.
We hence approach the problem via other methods, namely gradient Young measures with principal quasiconvexity. In fact, the gradient Young measures which the weakly convergent sequences generate often provide a convenient and flexible context for localisation arguments; we refer for the definitions and basic properties of Young measures to the monographs [39, 41, 42]. In particular, we will use the notation and terminology of gradient Young measures as described in our paper [4].
@@
The use of gradient Young measures also enables then new methods and views to analyse the localisation limits. In particular, it is here that the principal quasiconvexity shows its usefulness. For instance, in [3] one first observes that weakly converging sequences of quasiregular mappings generate gradient Young measures that are supported on the quasiconformal wells $Q_{2}(K)$ defined at (1.6). And second, thanks to Stoilow factorization, the properties of those Young measures can be analysed with help of principal maps, via the next
Theorem 1.3.
@@
Consider a gradient Young measure $\nu\in\mathscr{M}^{s}_{\textup{qc}}(Q_{2}(K))$, where $\frac{2K}{K+1}<s<\frac{2K}{K-1}$. Then there is a sequence of $K$-quasiconformal principal maps which generates in $\mathbb{D}$ the Young measure $\nu$.
@@
For mappings of bounded distortion as above, the classical Stoilow factorization is among the main tools in the above result. However, looking for similar methods for mappings of finite or integrable distortion, it is the Iwaniec–Šverák Stoilow factorization that provides us the following important improvement, see [4].
Theorem 1.4.
@@
Let $\nu\in\mathscr{M}^{2}_{\mathrm{qc}}(\mathbb{R}^{2\times 2}_{+})$ be generated by $(\textup{D}\psi_{j})$, where $(\psi_{j})$ is a sequence of homeomorphisms that is bounded in $\mathrm{W}^{1,2}(\mathbb{D})$ and such that $\int_{\mathbb{D}}\!K^{q}_{\psi_{j}}\,\mathrm{d}m(z)\leq C$ for some $q>1$. Then there is a sequence of maps $f_{j}\colon\mathbb{C}\to\mathbb{C}$ such that:
@@
    $(f_{j})\subset\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{D})$ is bounded and $(\textup{D}f_{j}|_{\mathbb{D}(0,R)})$ generates $\nu$ for all $R<1$;
@@
In order to apply these ideas we need that ${\bf E}(\textup{D}f_{j})$ is equi-integrable. Thus the proofs of two of our theorems split naturally into two propositions. The first states that the growth conditions, such as given in (1.11) below, guarantee that the sequence ${\bf E}(\textup{D}f_{j})$ is equi-integrable which rules out concentration effects. The second proposition, needed in Theorem 1.6, yields lower semicontinuity once equi-integrability is available, taking care of oscillation effects. The result is satisfactory but equi-integrability is a consequence of our assumption that the distortions are bounded in $\mathrm{L}^{1+\epsilon}$ for some $\epsilon>0$. As a matter of fact in [4] a similar issue arises from the Burkholder functional itself. Fortunately the scope of principal quasiconvexity and Stoilow factorization goes beyond oscillation effects and our third theorem shows that in the case of the Burkholder functional, principal quasiconvexity allows us to use a blow-up technique to prove lower semicontinuity in the borderline case, where equi-integrability cannot be assumed.
@@ 1.2. Main results
(1.8) $\mathbf{B}_{p}(A)\equiv\Bigl{(}(\tfrac{p}{2}-1)|A|^{2}-\tfrac{p}{2}\det A\Bigr{)}|A|^{p-2}.$
We note that $\mathbf{B}_{2}=-\det$ while for $p>2$, in terms of the $K$-quasiconformal well (1.6), we have $\mathbf{B}_{p}(A)\leq 0$ if and only if $A\in Q_{2}\left(\frac{p}{p-2}\right)$. Thus, fixing $K\equiv\frac{p}{p-2}$, or $p=\frac{2K}{K-1}$, one is led [4] to consider the local Burkholder functional
$\mathbf{B}_{K}(A)\equiv\left\{\begin{array}[]{ll}\mathbf{B}_{p}(A)&\mbox{ if }A\in Q_{2}(K)\\ +\infty&\mbox{ if }A\in\mathbb{R}^{2\times 2}\setminus Q_{2}(K).\end{array}\right.$
We emphasize the rather unusual feature of $\mathbf{B}_{K}$ that it is nonpositive and $p$-homogeneous on its effective domain $Q_{2}(K)$. As is well known, a planar $K$-quasiregular map lies in $\mathrm{W}^{1,p}_{\operatorname{loc}}$ for every $p<p_{K}\equiv\frac{2K}{K-1}$, while in general the integrability of the derivative fails at the borderline $p_{K}=\frac{2K}{K-1}$. Instead its derivative is locally in the Marcinkiewicz space $\mathrm{L}^{p_{K},\infty}$ (see [2]). On the other hand, a deep result of [6] states that for $K$-quasiregular maps $f\colon\Omega\to\mathbb{C}$, the integral
$0\geq\int_{C}\!{\bf B}_{K}(\textup{D}f)\,\mathrm{d}m(z)>-\infty$
for compact subsets $C$ of $\Omega$. However, even when $K$-quasiregular maps satisfy $f_{j}\rightharpoonup f$ weakly in $\mathrm{W}^{1,p_{K}}$ it still is far from clear whether $\bigl{(}\mathbf{B}_{p_{K}}(\textup{D}f_{j})\bigr{)}$ is locally equi-integrable. Thus the closed quasiconvexity of $\mathbf{B}_{K}$, proven in [4], led there to results that correspond to the $\mathrm{W}^{1,q}$-swlsc of $\mathbf{B}_{K}$ only for $q>p_{K}$. We refer to [19, 20] for the origins of the functional, to [6, 7, 9, 27, 44, 30] for its relation to quasiconvexity and to [5, 15, 16, 18, 40] for results in relation with Iwaniec’s conjecture.
In this note we in particular extend, using principal quasiconvexity and Stoilow factorization, the $\mathrm{W}^{1,q}$-swlsc of the local Burkholder functional up to the borderline case $q=p_{K}$. This is a consequence of a stronger and in a sense more natural result. To explain why the result is natural we fix a bounded open subset $\Omega$ of $\mathbb{C}$ and a $K$-quasiregular map $g\colon\mathbb{C}\to\mathbb{C}$, where $K>1$. Let $p=p_{K}\equiv 2K/(K-1)$ and consider the variational problem of minimizing the Burkholder energy
@@
over all $K$-quasiregular maps $v\in g+\mathrm{W}^{1,2}_{0}(\Omega)$. First one notes that from [6] (see also below) it follows that the Burkholder energy $\mathcal{B}[v]$ is bounded below over $v\in g+\mathrm{W}^{1,2}_{0}(\Omega)$. Next, let $(u_{j})$ be a minimizing sequence, so that in particular each $u_{j}\colon\Omega\to\mathbb{C}$ is $K$-quasiregular and, in view of the Dirichlet condition, the sequence $(u_{j})$ is bounded in $\mathrm{W}^{1,2}(\Omega)$. By a normality result for $K$-quasiregular maps [5], $(u_{j})$ admits a subsequence $(u_{j_{k}})$ that converges uniformly to a $K$-quasiregular map $u\in g+\mathrm{W}^{1,2}_{0}(\Omega)$. In order to conclude that $u$ is a minimizer one therefore needs lower semicontinuity of $\mathcal{B}[v]$ along the uniformly convergent subsequence $(u_{j_{k}})$. Apart from the technical assumption on the limit map $u$ the following result yields exactly that.
@@
Let $p>2$, $K\equiv p/(p-2)$, $\Omega$ be an open subset of $\mathbb{C}$ and suppose $u_{j}$, $u\colon\Omega\to\mathbb{C}$ are $K$-quasiregular maps such that $u_{j}\to u$ locally uniformly on $\Omega$. If $u\in\mathrm{W}^{1,p}_{\operatorname{loc}}(\Omega)$, then
@@
We refer to Corollary 3.7 for a global version of the Burkholder lower semicontinuity result on Dirichlet classes. While the result is stated for the Burkholder functional the interested reader will see that the proof also applies to the principal quasiconvex functionals of Theorem 2.2.
Note that in Theorem 1.5, since we do not know if the sequence $\mathbf{B}_{p}(\textup{D}u_{j})$ is locally equi-integrable, we need to handle the concentration and oscillation phenomena simultaneously. In this setting we will actually prove that the singular part of a suitably defined concentration measure (the reduced defect measure) is zero.
@@
The growth conditions we are able to deal with are the following,
(1.11) ${\bf E}^{-}(A)\leqslant C_{1}(\max\{|A|^{2-\epsilon},-\log(\det A),K_{A}\}+C_{2},$
where ${\bf E}^{-}(A)=\left|\min\{{\bf E}(A),0\}\right|$. With these bounds we have:
@@
Let ${\bf E}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ be lower semicontinuous, $\mathrm{W}^{1,1}$-principal quasiconvex and satisfying (1.11). Fix a homeomorphism $g\colon\mathbb{C}\to\mathbb{C}$ of class $\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{C})$ and assume it has integrable distortion on $\Omega$, $\|K_{g}\|_{\mathrm{L}^{1}(\Omega)}<+\infty$. Let $(u_{j})\subset g+\mathrm{W}^{1,2}_{0}(\Omega)$ with $\sup_{j}\|K_{u_{j}}\|_{\mathrm{L}^{q}(\Omega)}<\infty$ for some $q>1$ and suppose that $u_{j}\rightharpoonup u$ in $\mathrm{W}^{1,2}(\Omega)$, Then
@@
Our last result provides lower semicontinuity along sequences of maps that have asymptotically finite distortion.
@@
(1.12) $K(x)\equiv\limsup_{j\to\infty}\mathbf{K}(\textup{D}u_{j}(x))<+\infty\quad\mbox{ for a.e.}\;x\in\mathbb{A}.$
Our next result is a general lower semicontinuity result for sequences with asymptotically finite distortion. We emphasize that the mappings do not necessarily have integrable distortion and that the result can be localized at an arbitrary measurable set $\mathbb{A}$.
@@
Let $\mathbf{E}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ be Borel, lower semicontinuous on +$\mathbb{R}^{2\times 2}_{+}\cup\{0\}$, $\mathrm{W}^{1,2}$-principal quasiconvex and assume that
(1.13) $\bigl{|}\mathbf{E}(A)\bigr{|}\leq\mathbf{C}(K_{A})\bigl{(}|A|^{2}+1\bigr{)},\quad\forall A\in\mathbb{R}^{2\times 2}_{+}\cup\{0\}.$
Let $\mathbb{A}$ be a measurable subset of $\Omega$.

### v3

@@ 1.1. Principal maps
Next, we compare principal quasiconvexity with the classical notions of semiconvexity in the vectorial calculus of variations. Firstly, it is immediate from Jensen’s inequality for convex functions and (1.8) that any lower semicontinuous convex functional on $\mathbb{R}^{2\times 2}$ restricts to a $\mathrm{W}^{1,1}$-principal quasiconvex functional on $\mathbb{R}^{2\times 2}_{+}$. Next, notice that if $f\in\mathrm{C}^{1}(\mathbb{D})\cap\mathrm{W}^{1,\infty}(\mathbb{D})$ with $\det\textup{D}f>0$, and for some linear $A\in\mathbb{R}^{2\times 2}_{+}$ we have $f\in A+\mathrm{W}^{1,\infty}_{0}(\mathbb{D})$, so that $f$ is a classical test function for the $\mathrm{W}^{1,\infty}$-quasiconvexity inequality, then $f$ can be extended as principal map off the unit disc. Indeed, if $A(z)=b_{0}z+b_{1}\bar{z}$, then the map $f^{\rm prin}(z)\equiv f(z)\mathbf{1}_{\mathbb{D}}+\bigl{(}b_{0}z+b_{1}/z\bigr{)}\mathbf{1}_{\mathbb{C}\setminus\mathbb{D}}$ is principal in the sense of Definition 1.1. Therefore $\mathrm{W}^{1,\infty}$-principal quasiconvexity implies the standard quasiconvexity inequality for sufficiently smooth maps. A standard argument then yields that a $\mathrm{W}^{1,\infty}$-principal quasiconvex functional is locally rank-one convex on the interior of its effective domain and therefore that it is locally Lipschitz there (see [33]).
On the other hand, as we shall see in Section 2 below, the classical area formula implies that minus the determinant, $\det(A)$, is $\mathrm{W}^{1,1}$-principal quasiconvex (although not $\mathrm{W}^{1.p}$-quasiconvex for $1\leqslant p<2$) , while the determinant itself, $\det(A)$, is not. This simple example shows, in particular, that polyconvexity, and hence also quasiconvexity, does not imply principal quasiconvexity in general. For general functionals on $\mathbb{R}^{2\times 2}_{+}$ we present an additional local condition, see (2.11), which together with standard quasiconvexity implies principal quasiconvexity, see Proposition 2.6.
@@
In addition, the class of boundary values of all principal maps, i.e. of all the maps $f$ for which (1.11) applies, is much larger than just the linear ones. In fact, if $\Omega$ is a Jordan domain in $\overline{\mathbb{C}}$ with $\infty\in\Omega$, take a conformal map $\Psi\colon\mathbb{C}\setminus\overline{\mathbb{D}}\to\Omega$ with $\Psi(\infty)=\infty$. Then if the boundary $\partial\Omega$ is sufficiently regular, we can extend $\Psi$ to a principal map of $\overline{\mathbb{C}}$. In particular, we see that the set of boundary values of principal maps on $\partial\mathbb{D}$ coincides with the set of boundary values on $\partial\mathbb{D}$ of all such Riemann maps.
@@
Let us then informally explain why principal quasiconvexity is relevant for establishing lower semicontinuity of variational integrals. The proof of Morrey [66] for the lower semicontinuity of functionals with standard quasiconvex integrands goes basically as follows. Firstly, given a weakly converging sequence $(\psi_{j})$ of Sobolev functions, one localises the sequence to achieve convergence to a linear map. Second, one replaces this localized sequence with one of affine boundary values, $u_{j}-Az\in\mathrm{W}_{0}^{1,2}(\Omega)$, via a carefully chosen cutoff. And third, one applies directly the definition of quasiconvexity to each such $u_{j}$. However, such a strategy breaks down for functionals that are defined and finite only on $\mathbb{R}^{2\times 2}_{+}$, simply because the cut-off modification might give negative determinants on sets of positive measure.
@@
We must therefore approach the lower semicontinuity problem via other methods, for example gradient Young measures and the related closed quasiconvexity. However, to deal with gradient Young measures, and investigate whether quasiconvex functionals are closed quasiconvex in practise one needs information of the generating sequence of the Young measure. It is here that the principal quasiconvexity shows its usefulness. For instance, in [4] one first shows that weakly converging sequences of quasiregular mappings generate gradient Young measures that are supported on the quasiconformal wells $Q_{2}(K)$, as defined in (1.6). And second, thanks to Stoilow factorization, the properties of those Young measures can be analysed with the help of principal maps, via the next result from [5]. After these observations, the definition of principal quasiconvexity seems the natural one.
Proposition 1.3.
@@
Consider a homogeneous $\mathrm{W}^{1,s}$ gradient Young measure $\nu$ supported on the $K$-quasiconformal well $Q_{2}(K)$, where the exponent 2K/(K+1)<s<2K/(K1)2K/(K+1)<s<2K/(K-1). Then there is a sequence of $K$-quasiconformal principal maps which generates in $\mathbb{D}$ the Young measure $\nu$.
@@
For mappings of bounded distortion, aka the quasiconformal maps, the classical Stoilow factorization is among the main tools in the above result. However, looking for similar methods for mappings of finite or integrable distortion, it is the Iwaniec–Šverák version of Stoilow factorization that provides us the following important improvement, see [5, Theorem 4.4].
Proposition 1.4.
@@
Let $\nu$ be a homogeneous Young measure generated by $(\textup{D}\psi_{j})$, where $(\psi_{j})$ is a sequence of homeomorphisms that is bounded in $\mathrm{W}^{1,2}(\mathbb{D})$ and such that $\sup_{j}\int_{\mathbb{D}}\!K^{q}_{\psi_{j}}\,\mathrm{d}m(z)<+\infty$ for some $q>1$. Then there is a sequence of homeomorphisms $f_{j}\colon\mathbb{C}\to\mathbb{C}$ such that:
@@
    $(f_{j})$ is bounded in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{D})$ and $(\textup{D}f_{j})$ generates $\nu$;
@@
We recently proved [6] a more general version of this theorem that applies to sequences of maps $(\psi_{j})$ with distortions bounded merely in $\mathrm{L}^{1}$. In particular we emphasize that the maps $\psi_{j}$ there need not necessarily be homeomorphisms. The proof has some new ideas compared to [5] and will appear elsewhere.
In order to apply Theorem 1.4 for a functional $\bf E$}, we need that the sequence $\bigl{(}{\bf E}(\textup{D}f_{j})\bigr{)}$ is equi-integrable. Thus the proofs of results concerning such general functionals, considered in Theorems 1.6 and 1.9 with Corollary 1.7, split naturally into two propositions. The first states that the growth conditions, such as given in (1.17) below, together with a higher integrability of the distortion functions, guarantee that the sequence $\bigl{(}{\bf E}(\textup{D}f_{j})\bigr{)}$ is equi-integrable. This rules out concentration effects.
The second proposition, needed in Theorem 1.6, takes care of the oscillation effects and yields lower semicontinuity once equi-integrability is available. As a matter of fact, in [5] a similar issue arose for the important Burkholder functional itself. Fortunately the scope of principal quasiconvexity and Stoilow factorization goes beyond the basic oscillation effects. Indeed, our third result, Theorem 1.5 shows that in the case of the Burkholder functional, principal quasiconvexity allows us to use a blow-up technique to prove lower semicontinuity in the borderline case, where equi-integrability cannot be assumed.
@@ 1.2. Main results
(1.13) $\mathbf{B}_{p}(A)\equiv\Biggl{(}\bigl{(}\frac{p}{2}-1\bigr{)}|A|^{2}-\frac{p}{2}\det A\Biggr{)}|A|^{p-2},\qquad A\in\mathbb{R}^{2\times 2}.$
We note that $\mathbf{B}_{2}(A)=-\det(A)$, while for $p>2$, in terms of the quasiconformal well (1.6), we have $\mathbf{B}_{p}(A)\leq 0$ if and only if $A\in Q_{2}\bigl{(}p/(p-2)\bigr{)}$. Thus, fixing $K\equiv p/(p-2)$, or equivalently, p=2K/(K1)p=2K/(K-1), one is led [5] to consider the local Burkholder functional defined as
(1.14) $\mathbf{B}_{K}^{\operatorname{loc}}(A)\equiv\left\{\begin{array}[]{ll}\mathbf{B}_{p}(A)&\mbox{ if }A\in Q_{2}(K)\\ +\infty&\mbox{ if }A\in\mathbb{R}^{2\times 2}\setminus Q_{2}(K).\end{array}\right$.
For the origins of the Burkholder functional we refer to [26, 27]. Its relation to Morrey’s problem about rank-one and quasiconvexity was first observed by T. Iwaniec and it has since been the subject of intense investigation by a number of authors, including [10, 11, 13, 45, 46, 81, 52]. See also [8, 21, 22, 25, 70, 72] for results about the norm of the Beurling-Ahlfors operator and the Iwaniec conjecture. We emphasize the rather unusual feature of $\mathbf{B}_{K}^{\operatorname{loc}}$ that it is nonpositive and $p$-homogeneous on its effective domain $Q_{2}(K)$.
As is well known, a planar $K$-quasiregular map lies in $\mathrm{W}^{1,p}_{\operatorname{loc}}$ for every $p<p_{K}\equiv 2K/(K-1)$, while in general the integrability of the derivative of such a map fails at the borderline exponent $p_{K}$. Instead, in general its derivative is locally in the Marcinkiewicz space $\mathrm{L}^{p_{K},\infty}$ (see [3]). On the other hand, an interesting result of [10] states that for $K$-quasiregular maps $f\colon\Omega\to\mathbb{C}$ the function $\mathbf{B}_{p_{K}}(\textup{D}f)$ is locally integrable on $\Omega$, that is,
$0\geq\int_{C}\!\mathbf{B}_{p_{K}}(\textup{D}f)\,\mathrm{d}m(z)>$-$\infty$
holds for compact subsets $C$ of $\Omega$. However, even when $K$-quasiregular maps satisfy $f_{j}\rightharpoonup f$ weakly in $\mathrm{W}^{1,p_{K}}$ it still is not clear whether $\bigl{(}\mathbf{B}_{p_{K}}(\textup{D}f_{j})\bigr{)}$ is locally equi-integrable. Thus the closed quasiconvexity of $\mathbf{B}_{K}^{\operatorname{loc}}$, proven in [5, Theorem 1.4], led there to results that correspond to the $\mathrm{W}^{1,q}$-swlsc of $\mathbf{B}_{K}^{\operatorname{loc}}$ only for $q>p_{K}$.
In this note we in particular extend, using principal quasiconvexity and Stoilow factorization, the $\mathrm{W}^{1,q}$-swlsc of the local Burkholder functional up to the borderline case $q=p_{K}$. This is a consequence of a stronger and in a sense more natural result. To explain why our result is natural we fix a bounded domain $\Omega$ of $\mathbb{C}$ and a $K$-quasiregular map $g\colon\mathbb{C}\to\mathbb{C}$, where $K>1$. Let $p=p_{K}$ be the corresponding borderline exponent and consider the variational problem of minimizing the Burkholder energy
@@
over all $K$-quasiregular maps $v\in g+\mathrm{W}^{1,2}_{0}(\Omega)$. First one notes that from [10] (see also Proposition 3.6 below for a more precise lower bound) it follows that the Burkholder energy $\mathcal{B}[v]$ is bounded below over $v\in g+\mathrm{W}^{1,2}_{0}(\Omega)$. Next, let $(u_{j})\subset g+\mathrm{W}^{1,2}_{0}(\Omega)$ be a minimizing sequence, so that in particular each $u_{j}\colon\Omega\to\mathbb{C}$ is $K$-quasiregular and, in view of the Dirichlet condition, the sequence $(u_{j})$ is bounded in $\mathrm{W}^{1,2}(\Omega)$. By a normality result for $K$-quasiregular maps [8], $(u_{j})$ admits a subsequence $(u_{j_{k}})$ that converges uniformly to a $K$-quasiregular map $u\in g+\mathrm{W}^{1,2}_{0}(\Omega)$. In order to conclude that $u$ is a minimizer one therefore needs lower semicontinuity of $\mathcal{B}[v]$ along the uniformly convergent subsequence $(u_{j_{k}})$. Apart from the technical assumption on the limit map $u$ the following result yields exactly that.
@@
Let $p>2$, $K\equiv p/(p-2)$, $\Omega$ be a domain in $\mathbb{C}$ and suppose $u_{j}$, $u\colon\Omega\to\mathbb{C}$ are $K$-quasiregular maps such that $u_{j}\to u$ locally uniformly on $\Omega$. If $u\in\mathrm{W}^{1,p}_{\operatorname{loc}}(\Omega)$, then
@@
Since any locally uniform limit $u$ of a sequence $(u_{j})$ of $K$-quasiregular maps is itself $K$-quasiregular, we see from [3] that the limit function $u$ in Theorem 1.5 automatically has the Marcinkiewicz-regularity $\textup{D}u\in\mathrm{L}^{p,\infty}_{\operatorname{loc}}$. It remains open if merely this suffices in Theorem 1.5.
We refer to Corollary 3.7 for a global version of the Burkholder lower semicontinuity result on Dirichlet classes. Since we do not know if the sequence $\bigl{(}\mathbf{B}_{p}(\textup{D}u_{j})\bigr{)}$ in Theorem 1.5 is locally equi-integrable, we need to handle the concentration and oscillation phenomena simultaneously. In this setting we will actually prove that the singular part of the reduced defect measure for the sequence $\bigl{(}\mathbf{B}_{p}(\textup{D}u_{j})\bigr{)}$ is zero.
@@
For simplicity we restrict attention here to energy densities satisfying the following growth conditions: There exist $C>0$ and $p\in[1,2)$ such that
(1.17) $\left\{\begin{array}[]{l}\bigl{|}{\bf E}(A)\bigr{|}\leq C\max\bigl{\{}|A|^{p},-\log(\det A),K_{A}\bigr{\}}+C\,\mbox{ on }\,\mathbb{R}^{2\times 2}_{+},\\ {}\\ {\bf E}(A)=+\infty\,\mbox{ on }\,\mathbb{R}^{2\times 2}\setminus\bigl{(}\mathbb{R}^{2\times 2}_{+}\cup\{0\}\bigr{)}.\end{array}\right$.
Note we do not impose any condition for the value of $\bf E$} at 0 and that the functional $\bf W$} defined at (2.5) satisfies (1.17) with $p=1$. With these bounds we have:
@@
Let ${\bf E}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ be Borel measurable and satisfy (1.17). Assume that $\bf E$} is $\mathrm{W}^{1,1}$ principal quasiconvex. Fix an exponent $q>\frac{p}{2-p}$ and let $g\colon\mathbb{C}\to\mathbb{C}$ be a homeomorphism of class $\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{C})$ with distortion $K_{g}\in\mathrm{L}^{q}_{\operatorname{loc}}(\mathbb{C})$.
Let $(u_{j})$ be a sequence of homeomorphisms in $g+\mathrm{W}^{1,2}_{0}(\Omega)$ with $\sup_{j}\|K_{u_{j}}\|_{\mathrm{L}^{q}(\Omega)}<\infty$ and suppose that $u_{j}\rightharpoonup u$ in $\mathrm{W}^{1,2}(\Omega)$. Then $u\in g+\mathrm{W}^{1,2}_{0}(\Omega)$ is a homeomorphism with $K_{u}\in\mathrm{L}^{q}(\Omega)$ and
@@
This result yields in a standard way an existence result for a variational problem related to hyperelasticity:
Corollary 1.7.
Let ${\bf E}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ be Borel measurable, satisfy (1.17) and be $\mathrm{W}^{1,1}$-principal quasiconvex. Suppose ${\bf P}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ is lower semicontinuous, polyconvex and satisfies the coercivity condition
(1.18) ${\bf P}(A)\geq c\bigl{(}|A|^{2}+K_{A}^{q}\bigr{)}$
for all $A\in\mathbb{R}^{2\times 2}$, where $c>0$, $q>\frac{p}{2-p}$ are constants. For a homeomorphism $g\colon\mathbb{C}\to\mathbb{C}$ of class $\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{C})$ and with ${\bf E}(\textup{D}g)+{\bf P}(\textup{D}g)\in\mathrm{L}^{1}(\Omega)$ we put
[…]
Then, for each $F\in\mathrm{W}^{-1,2}(\Omega)\equiv\mathrm{W}^{1,2}_{0}(\Omega)^{\ast}$, the variational problem
$\inf_{u\in\mathcal{A}_{g}}\left(\int_{\Omega}\!\bigl{(}{\bf E}(\textup{D}u)+{\bf P}(\textup{D}u)\bigr{)}\,\mathrm{d}m(z)+\bigl{\langle}F,u\bigr{\rangle}\right)$
admits a minimizer $u\in\mathcal{A}_{g}$ with $K_{u}\in\mathrm{L}^{q}(\Omega)$.
Our last result provides lower semicontinuity for principal quasiconvex functionals along sequences of maps that have asymptotically finite distortion in the following sense.
@@
(1.19) $\limsup_{j\to\infty}\mathbf{K}(\textup{D}u_{j}(x))<+\infty\quad\mbox{ for a.e.}\;x\in\mathbb{A}.$
We emphasize that the mappings do not necessarily have finite distortion everywhere, but that the distortion condition is merely required to hold on the measurable set $\mathbb{A}$.
@@
Let $\mathbf{E}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ be lower semicontinuous and $\mathrm{W}^{1,2}$-principal quasiconvex, with $\bf E$$\infty$ on $\mathbb{R}^{2\times 2}\setminus\bigl{(}\mathbb{R}^{2\times 2}_{+}\cup\{0\}\bigr{)}$, and assume that
(1.20) $\bigl{|}\mathbf{E}(A)\bigr{|}\leq\mathbf{C}(K_{A})\bigl{(}|A|^{2}+1\bigr{)},\quad\forall A\in\mathbb{R}^{2\times 2}_{+},$
where $\mathbf{C}\colon[1,\infty)\to[1,\infty)$ is an increasing function. Let $\mathbb{A}$ be a measurable subset of $\Omega$.